### v1

@@ 1 Introduction
Recovering a joint distribution from its marginal distributions is a fundamental problem in statistics, economics, machine learning, and related fields Sklar (1959); Genest & Rivest (1993); Li (1999); Beare (2010); Joe (2014); McNeil et al. (2015); Fan & Henry (2023); Chambers et al. (2025). However, marginal distributions alone are generally insufficient to uniquely determine the joint distribution–infinitely many joint distributions may be consistent with the same set of marginals. This ambiguity grows with dimensionality, as the number of free parameters in the joint distribution increases exponentially with the number of random variables in discrete statistical models.
To resolve this indeterminacy, additional structure must be imposed–such as assumptions about dependence patterns, joint moments, or other inter-variable relationships. A fundamental approach is to introduce a functional mapping, denoted by $f$, that takes a set of discrete marginal distributions and returns a unique joint distribution consistent with them. We refer to such a mapping $f$ as a model. Each model imposes certain structural assumptions, such as positive dependence, negative dependence, and independence.
@@
To illustrate, consider electricity and water consumption over a fixed period of time in a neighborhood. Each consumer in the neighborhood consumes some amount of water, and some amount of electricity (in the jargon of economists, we speak of them choosing water consumption and electricity consumption). Water and electricity consumption are each naturally modeled as real-valued, but we suppose that each of these these real-valued consumption levels are partitioned into bins: low, medium and high: ($L$, $M$, and $H$). Each bin represents an interval of possible real-valued consumption levels. Importantly, a consumer whose water choice lies in bin $L$ consumes less water than one whose choice lies in bin $M$; we write $L<M<H$. The economist observes the empirical distribution of choices from bins $L$, $M$ and $H$ for both water consumption across all consumers, and for electricity consumption. We refer to these as marginals.
Even though each consumer chooses both water and electricity consumption, we assume that these joint choices are unobserved. That is, we imagine data comes from different sources, leaving the joint distribution of water and electricity consumption unknown. We seek to study methods of selecting joint distributions. As a modeler or policymaker, we may hypothesize a positive association–for instance, we may conjecture that higher electricity usage may tend to coincide with higher water usage. Given the marginals, our model produces a corresponding joint distribution. But many other hypotheses are possible.
@@
Violating IA would render social policies sensitive to arbitrary data-reporting choices–especially problematic when reporting methods vary across contexts. Worse, it could allow unscrupulous data collectors to manipulate outcomes by altering data presentation. Many, but not all, models satisfy IA. A trivial example of a model satisfying IA is the independence model, which assumes statistical independence across variables, constructing the joint distribution via multiplication of marginals. However, this is not the only such model. Our main result provides a complete characterization of all IA-compliant models: they are precisely those defined by copulas, a well-studied class of statistical objects. Indeed, for statisticians, to get from marginal distributions on the real line to a joint distribution, copulas come to mind first Nelsen (2006). Sklar’s theorem states that a joint distribution on an Euclidean space can be described by its marginal distributions and a copula. Our task is, however, different: in our setting, both the marginal distributions and the joint distribution vary (and they are categorical), while keeping the invariance property.
@@
Other sections of the paper are organized as follows. Section 1.1 introduces the notation and definitions of IA. Section 1.2 introduces and proves the main result of the paper. Section 1.3 investigated an additional result characterizing a family of independence copulas. Finally, Section 1.4 concludes.
1.1 The Model
@@ 44 $1/32$ $1/32$ $1/32$ $1/32$ $1/8$
Take any $w\in\{2,3,\dots\}$ and $q\in\Delta([w])$. For $j\in[w-1]$, let $q^{j}\in\Delta([w-1])$ be given by, for all $t\in[w-1]$,
$q^{j}(t)=q(t)\mathds{1}_{\{t\leq j\}}+q(t+1)\mathds{1}_{\{t\geq j\}}.$
@@ 11 $1/4$ 0 $1/8$ $1/8$ $1/2$
There are many copulas, a good introduction to the theory is Nelsen (2006). Two classical ones are as follows:
@@
for all $p\in\Delta^{n}_{\rm ma}$ and $s\in\mathbb{R}^{n}$. The left hand side of (2) reports the cumulative probability $f(p)$ assigns to $s$, that is, the probability of $\{s^{\prime}:s^{\prime}_{i}\leq s_{i}\mbox{ for all }i\}$. As in the standard single-dimensional case, this object uniquely pins down $f(p)$, through an inclusion-exclusion principle. The right hand side applies the copula $C$ to the cumulative probabilites of the marginals $p_{i}$.
If $C$ is the independence copula, the corresponding copula model is the independence model. The Fréchet-Hoeffding upper bound model is closely related to the progressive random choice model of Filiz-Ozbay & Masatlioglu (2023).
We now show what happens with the application of the Fréchet-Hoeffding upper bound. First, consider $(p_{1},p_{2})$, then the Fréchet-Hoeffding upper bound model results in
@@ 1.2 Main Result
  • Our notion of a model allows any $z\in\mathbb{N}^{N}$. However, we could restrict the domain of $z$ to any $\prod_{i}[1,\hat{z}_{i}]$, where $\hat{z}_{i}\in\mathbb{N}\cup\{+\infty\}$. So long as $\hat{z}_{i}\geq 3$, the characterization still holds (a different, more tedious proof is necessary, however).
@@
1.3 Unordered categorical variables
@@
Inspired by this observation, we propose the following symmetry property that removes the location from points in $[w]$. Let $\Pi_{w}$ be the set of all permutations on $[w]$ for $w\in\mathbb{N}$. For $q\in\Delta([w])$ and $\sigma\in\Pi_{w}$, $q^{\sigma}(s)=q(\sigma(s))$ for all $s\in[w]$. Let $T_{i}^{\sigma}$ be the mapping that maps $p$ to $(p_{1},\dots,p_{i-1},p_{i}^{\sigma},p_{i+1}\dots,p_{n})$, for all $i\in[n]$ and $\sigma\in\Pi_{z_{i}}$. When $T_{i}^{\sigma}$ is applied to a vector $s$, it means $T_{i}^{\sigma}(s)=(s_{1},\dots,s_{i-1},\sigma(s_{i}),s_{i+1},\dots,s_{n})$.
@@
  1. (i)
    Neutrality: For $p\in\Delta_{\rm ma}^{n}$, $i\in N$, $\sigma\in\Pi_{z_{i}}$ and $s\in[z]$, $f(T^{\sigma}_{i}(p))(s)=f(p)(T^{\sigma}_{i}(s)).$
  2. (ii)
    $M$-neutrality: For $p\in\Delta_{\rm ma}^{n}$, $i\in M$, $\sigma\in\Pi_{z_{i}}$ and $s\in[z]$, $f(T^{\sigma}_{i}(p))(s)=f(p)(T^{\sigma}_{i}(s)).$
@@
for all $\sigma\in\Pi_{k}$. Repeatedly applying this equality, we get that, for each given $(s_{i})_{i\in N\setminus M}$, $P_{C}$ assigns equal probability to each point in the set $\{(s_{1}/k,\dots,s_{n}/k):(s_{i})_{i\in M}\in[k]^{m}\}$. Since $k$ is arbitrary, we know that the $M$-marginal distribution of $P_{C}$ is uniform on $0,1]^{m}$, and under $P_{C}$, the components in $M$ are independent of all remaining components in $N\setminus M$. Hence, the copula has the form $C(x)=(\prod_{i\in M}x_{i})\tilde{C}((x_{j})_{j\in N\setminus M})$.
@@
1.4 Conclusion
In this paper, we proposed and characterized the property of Invariant Aggregation (IA)–a condition that requires merging adjacent categories in marginal distributions not to distort the joint distribution over unaffected categories. We proved that IA holds if and only if the model is based on a copula, thereby offering a precise, axiomatic foundation for when and why copula-based modeling is appropriate in discrete choice settings. This result highlights the powerful structure of copulas in maintaining consistency under data coarsening or refinement, which is common in empirical contexts.
@@ 1.4 Conclusion
Recent contributions focus on ordered choices. A major contribution is due to Apesteguia et al. (2017); Filiz-Ozbay & Masatlioglu (2023), the so-called progressive random choice (PRC) model (see also Yildiz (2022); Apesteguia & Ballester (2023); Petri (2023); Masatlioglu & Vu (2024)). In this model, the set $X$ is endowed with a linear order and each $A_{i}\subseteq X$ inherits the restriction of this order. The progressive random choice model coincides with the Fréchet-Hoeffding upper bound model discussed here Chambers et al. (2025). In economics problems, we sometimes imagine new data arriving in the form of a new subset $B\subseteq X$ and a distribution over $B$. We realized that the application of the PRC model when including $B$ in the data did not distort the marginal probabilities assigned to each member of $\bigtimes_{i}A_{i}$. This naturally led to the investigation of the IA property here.

### v2

@@ 1 Introduction
Recovering a joint distribution from its marginal distributions is a fundamental problem in statistics, economics, machine learning, and related fields Sklar (1959); Genest & Rivest (1993); Li (1999); Beare (2010); Joe (2014); McNeil et al. (2015); Fan & Henry (2023); Chambers et al. (2025). However, marginal distributions alone are generally insufficient to uniquely determine the joint distribution—infinitely many joint distributions may be consistent with the same set of marginals. This ambiguity grows with dimensionality, as the number of free parameters in the joint distribution increases exponentially with the number of random variables in discrete statistical models.
To resolve this indeterminacy, additional structure must be imposed—such as assumptions about dependence patterns, joint moments, or other inter-variable relationships. A fundamental approach is to introduce a functional mapping, denoted by $f$, that takes a set of discrete marginal distributions and returns a unique joint distribution consistent with them. We refer to such a mapping $f$ as a model. Each model imposes certain structural assumptions, such as positive dependence, negative dependence, and independence.
@@
To illustrate, consider electricity and water consumption over a fixed period of time in a neighborhood. Each consumer in the neighborhood consumes some amount of water, and some amount of electricity (in the jargon of economists, we speak of them choosing water consumption and electricity consumption). Water and electricity consumption are each naturally modeled as real-valued, but we suppose that each of these real-valued consumption levels is partitioned into bins: low, medium, and high: ($L$, $M$, and $H$). Each bin represents an interval of possible real-valued consumption levels. Importantly, a consumer whose water choice lies in bin $L$ consumes less water than one whose choice lies in bin $M$; we write $L<M<H$. The economist observes the empirical distribution of choices from bins $L$, $M$ and $H$ for both water consumption across all consumers, and for electricity consumption. We refer to these as marginals.
Even though each consumer chooses both water and electricity consumption, we assume that these joint choices are unobserved. That is, we imagine data comes from different sources, leaving the joint distribution of water and electricity consumption unknown. We seek to study methods of selecting joint distributions. As a modeler or policymaker, we may hypothesize a positive association—for instance, we may conjecture that higher electricity usage may tend to coincide with higher water usage. Given the marginals, our model produces a corresponding joint distribution. But many other hypotheses are possible.
@@
Violating IA would render social policies sensitive to arbitrary data-reporting choices—especially problematic when reporting methods vary across contexts. Worse, it could allow unscrupulous data collectors to manipulate outcomes by altering data presentation. Many, but not all, models satisfy IA. A trivial example of a model satisfying IA is the independence model, which assumes statistical independence across variables, constructing the joint distribution via multiplication of marginals. However, this is not the only such model. Our main result provides a complete characterization of all IA-compliant models: they are precisely those defined by copulas, a well-studied class of statistical objects. Indeed, for statisticians, to get from marginal distributions on the real line to a joint distribution, copulas come to mind first Nelsen (2006). Sklar’s theorem states that a joint distribution on an Euclidean space can be described by its marginal distributions and a copula. Our task is, however, different: in our setting, both the marginal distributions and the joint distribution vary (and they are categorical), while keeping the invariance property.
@@
Other sections of the paper are organized as follows. Section 2 introduces the notation and definitions of IA. Section 3 introduces and proves the main result of the paper. Section 4 investigated an additional result characterizing a family of independence copulas. Finally, Section 5 concludes.
2 The Model
@@ 44 $1/32$ $1/32$ $1/32$ $1/32$ $1/8$
Take any $w\in\{2,3,\dots\}$ and $q\in\Delta([w])$. For $j\in[w-1]$, let $q^{j}\in\Delta([w-1])$ be given by, for all $t\in[w-1]$, $q^{j}(t)=q(t)\mathds{1}_{\{t\leq j\}}+q(t+1)\mathds{1}_{\{t\geq j\}}.$
@@ 11 $1/4$ 0 $1/8$ $1/8$ $1/2$
There are many copulas; a good introduction to the theory is Nelsen (2006). [^3] Copulas have also been applied in other areas of economic theory, such as bargaining Bastianello & LiCalzi (2019) and behavioral game theory Frick et al. (2022). Two classical ones are as follows:
@@
for all $p\in\Delta^{n}_{\rm ma}$ and $s\in\mathbb{R}^{n}$. The left-hand side of (2) reports the cumulative probability $f(p)$ assigned to $s$, that is, the probability of $\{s^{\prime}:s^{\prime}_{i}\leq s_{i}\mbox{ for all }i\}$. As in the standard single-dimensional case, this object uniquely pins down $f(p)$, through an inclusion-exclusion principle. The right-hand side applies the copula $C$ to the cumulative probabilities of the marginals $p_{i}$.
If $C$ is the independence copula, the corresponding copula model is the independence model. We now demonstrate what happens when the Fréchet-Hoeffding upper bound is applied. First, consider $(p_{1},p_{2})$, then the Fréchet-Hoeffding upper bound model results in
@@ 3 Main Result
  • Our notion of a model allows any $z\in\mathbb{N}^{N}$. However, we could restrict the domain of $z$ to any $\times_{i}[1,\hat{z}_{i}]$, where $\hat{z}_{i}\in\mathbb{N}\cup\{+\infty\}$. So long as $\hat{z}_{i}\geq 3$, the characterization still holds (a different, more tedious proof is necessary, however).
@@
4 Unordered categorical variables
@@
Inspired by this observation, we propose the following symmetry property that removes the location from points in $[w]$. Let $\Sigma_{w}$ be the set of all permutations on $[w]$ for $w\in\mathbb{N}$. For $q\in\Delta([w])$ and $\sigma\in\Sigma_{w}$, $q^{\sigma}(s)=q(\sigma(s))$ for all $s\in[w]$. Let $T_{i}^{\sigma}$ be the mapping that maps $p$ to $(p_{1},\dots,p_{i-1},p_{i}^{\sigma},p_{i+1}\dots,p_{n})$, for all $i\in[n]$ and $\sigma\in\Sigma_{z_{i}}$. When $T_{i}^{\sigma}$ is applied to a vector $s$, it means $T_{i}^{\sigma}(s)=(s_{1},\dots,s_{i-1},\sigma(s_{i}),s_{i+1},\dots,s_{n})$.
@@
  1. $M$-neutrality: For $p\in\Delta_{\rm ma}^{n}$, $i\in M$, $\sigma\in\Sigma_{z_{i}}$ and $s\in[z]$, $f(T^{\sigma}_{i}(p))(s)=f(p)(T^{\sigma}_{i}(s)).$
@@
for all $\sigma\in\Sigma_{k}$. Repeatedly applying this equality, we get that, for each given $(s_{i})_{i\in N\setminus M}$, $P_{C}$ assigns equal probability to each point in the set $\{(s_{1}/k,\dots,s_{n}/k):(s_{i})_{i\in M}\in[k]^{m}\}$. Since $k$ is arbitrary, we know that the $M$-marginal distribution of $P_{C}$ is uniform on $0,1]^{m}$, and under $P_{C}$, the components in $M$ are independent of all remaining components in $N\setminus M$. Hence, the copula has the form $C(x)=(\prod_{i\in M}x_{i})\tilde{C}((x_{j})_{j\in N\setminus M})$.
@@
5 Conclusion
In this paper, we proposed and characterized the property of Invariant Aggregation (IA)—a condition that requires merging adjacent categories in marginal distributions not to distort the joint distribution over unaffected categories. We proved that IA holds if and only if the model is based on a copula, thereby offering a precise, axiomatic foundation for when and why copula-based modeling is appropriate in discrete choice settings. This result highlights the powerful structure of copulas in maintaining consistency under data coarsening or refinement, which is common in empirical contexts.
@@ 5 Conclusion
Recent contributions focus on ordered choices. A major contribution is due to Apesteguia et al. (2017); Filiz-Ozbay & Masatlioglu (2023), the so-called progressive random choice (PRC) model (see also Yildiz (2022); Apesteguia & Ballester (2023); Petri (2023); Masatlioglu & Vu (2024)). In this model, the set $X$ is endowed with a linear order and each $A_{i}\subseteq X$ inherits the restriction of this order. Chambers et al. (2025) introduces copula-based models for stochastic choice and shows that the progressive random choice model is a special case. Indeed, the progressive random choice model coincides with the Fréchet-Hoeffding upper bound model. In economics problems, we sometimes imagine new data arriving in the form of a new subset $B\subseteq X$ and a distribution over $B$. We realized that the application of the PRC model when including $B$ in the data did not distort the marginal probabilities assigned to each member of $\bigtimes_{i}A_{i}$. This naturally led to the investigation of the IA property here.